7. A ball is thrown vertically upward from the ground with an initial speed of 15.0m/s.
a) How long does it take the ball to reach its maximum altitude?
b) What is its maximum altitude?
c) How long did the ball stayed in air.
d) Determine the velocity of the ball at t= 2.00s.
1
Expert's answer
2020-05-03T17:17:55-0400

(a)

"t=\frac{v_0}{g}=\frac{15.0}{9.81}=1.53\:\rm s."

(b)

"h_{\rm max}=\frac{v_0^2}{2g}=\frac{15.0^2}{2\times 9.81}=11.5\:\rm m."

(c)

"\tau=2t=2\times 1.53=3.06\:\rm s."

(d)

"v=v_0-gt=15.0-9.81\times 2.00=-4.62\:\rm m/s"